An electric bulb of volume 250 cm^3 was sealed off during manufacture at a pressure of 10^-3 mm Hg at a 27 degree celsius,find the number of molecules in it.(Avogadro number=6*10^23)

Grade:11

2 Answers

Arun Kumar IIT Delhi
9 years ago

Now apply PV=nRT


Get n and multiply with Avogadro's Constant.


Arun Kumar
IIT Delhi
Askiitians Faculty
Akshita
5 years ago
PV=k×N×T where we have to find N that is number of molecules. Given V=250 cm3 , P=10-3 mm Hg , T= 27°C